OpenStudy (anonymous):

Find the limit as x approaches infinity and negative infinity. Use that info, together with intercepts, to give a rough sketch of the graph y=(1-x)^2(2-x)(5-x)

OpenStudy (anonymous):



OpenStudy (rational):

1, 2, 5 are your intercepts right ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

why did u mark "-1" in above diagram ?

OpenStudy (anonymous):

1-x^2=0 1=x^2 =+/-1?

OpenStudy (rational):

Nope. what you're given is (1-x)^2 and not 1-x^2

OpenStudy (anonymous):

Oh oops!

OpenStudy (rational):